Mathematics 134 – Intensive Calculus for Science 2 Quiz 1 – January 27, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes and you may use the calculator anywhere you like.

A) (10) The following graph shows y = f(x). Use it to determine ∫_{0}^{3} f(x) dx. (Hint: think area)

B) (5) What is the average value of f(x) in question A for 0 ≤ x ≤ 3.

C) (10) Use the Fundamental Theorem of Calculus to evaluate ∫_{1}^{2} x + \sqrt{x} dx.